OpenStudy (anonymous):

what is the standard deviation for the numbers 9,10,10,8,7,11,12,9

OpenStudy (anonymous):

find the mean then subtract the mean from each other numbers

OpenStudy (anonymous):

then square each one

OpenStudy (anonymous):

add all those numbers up then divide that number by how many values their are then square that answer then tada

OpenStudy (kirbykirby):

Assuming this is a sample: \[ \large s=\sqrt{\frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2}\]

OpenStudy (anonymous):

u wanna do it that way u can

OpenStudy (anonymous):

hmm i think i did something wrong with my calculations. I'm confused now

OpenStudy (anonymous):

it tells me to round to two decimal places